Suppose that each time Bob shoots a free throw, he has a 3/4 probability of success. If Bob shoots three free throws, what is the probability that he succeeds on at least two of them?
Answer #1

X ~ Bin ( n , p)

Where n = 3 , p = 3/4

Binomial probability distribution is

P(X) = nCx px ( 1 - p)n-x

So,

P(X >= 2) = P(X = 2) + (X = 3)

= 3C2* ( 3/4)2 * ( 1 - 3/4) + 3C3* ( 3/4)3 * ( 1 - 3/4)0

= 0.8438